OpenStudy (anonymous):

solve this system of conic sections: x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0

jimthompson5910 (jim_thompson5910):

Subtract the second equation from the first to get x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0 ------------------- 0x^2 + 0y^2 - 2x - 4y - 12 = 0 So we now have the equation -2x - 4y - 12 = 0 Now solve for x to get -2x - 4y - 12 = 0 -2x - 4y = 12 -2x = 12 + 4y x = (12 + 4y)/(-2) x = -6-2y

jimthompson5910 (jim_thompson5910):

Now plug this into any equation. Let's pick the first one x^2+y^2+2x+2y=0 (-6-2y)^2+y^2+2(-6-2y)+2y=0 ... plug in x = -6 - 2y (-6-2y)(-6-2y)+y^2+2(-6-2y)+2y=0 -6(-6-2y)-2y(-6-2y)+y^2+2(-6-2y)+2y=0 36+12y+12y+4y^2+y^2-12-4y+2y=0 5y^2+22y+24=0 (y+2)(5y+12)=0 y+2 = 0 or 5y+12 = 0 y = -2 or 5y = -12 y = -2 or y = -12/5 Now use those two y values above to find the corresponding x values. Tell me what you get.

OpenStudy (anonymous):

ok hold on while i solve

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

ok well for when y=-2 i get x=0 or x=-2 now for when y=-12/5 i am stuck this is what is get

jimthompson5910 (jim_thompson5910):

plug those y values into x = -6 - 2y

OpenStudy (anonymous):

oh ok oops ok hold on a sec

OpenStudy (anonymous):

ok i get x= -2 and x= -1.2

jimthompson5910 (jim_thompson5910):

-1.2 = -6/5 so the two ordered pair solutions are (-2,-2) and (-6/5,-12/5)

OpenStudy (anonymous):

why does -1.2 = -6/5?

jimthompson5910 (jim_thompson5910):

-1.2 = -(1.2) -1.2 = -(1+ 0.2) -1.2 = -(1+ 20/100) -1.2 = -(100/100+ 20/100) -1.2 = -120/100 -1.2 = -6/5

OpenStudy (anonymous):

oh duh ok i get it lol thanks so much (: can you check two or three problems i did today to see if i did them right plz?

jimthompson5910 (jim_thompson5910):

Also, if you plug in y = -12/5 and keep everything as a fraction, you get this... x = -6-2y x = -6-2(-12/5) x = -6+24/5 x = -30/5+24/5 x = (-30 + 24)/5 x = -6/5

jimthompson5910 (jim_thompson5910):

sure go ahead and post them